When the vertical lifting of a body weighing 2 kg to a height of 10 m was performed work 240 J. Determine the acceleration of the body. 


1
Expert's answer
2021-07-21T14:41:45-0400

Work (w)=240J

Mass(m)=2kg

Height (s)=10m

Work=F"\times" s

"F=\frac{w}{s}"

"F=\frac{240}{10}=24N"

We know that

F=ma

"a=\frac{F}{m}=\frac{24}{2}=12m/sec^2"